27 drops of smalll size are charged at 220v each.They coalesce to form a bigger drop.calculate the potential of the bigger drop?

Grade:

1 Answers

Chetan Mandayam Nayakar
12 years ago

Capacitance C=4πε0r, let radius of big drop be R,R=3r, Q=27q,V=Q/4πε0R=27*220/3

Dear moderator, please email me at mn_chetan@yahoo.com, in case you feel solution given above is insufficient